### v1

# Parameter-Efficient Fine-Tuning via Circular Convolution

Abstract
Low-Rank Adaptation (LoRA) has gained popularity for fine-tuning large foundation models, leveraging low-rank matrices $\mathbf{A}$ and $\mathbf{B}$ to represent weight changes (i.e., $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$). This method reduces trainable parameters and mitigates heavy memory consumption associated with full delta matrices by sequentially multiplying $\mathbf{A}$ and $\mathbf{B}$ with the activation. Despite its success, the intrinsic low-rank characteristic may limit its performance. Although several variants have been proposed to address this issue, they often overlook the crucial computational and memory efficiency brought by LoRA. In this paper, we propose Circular Convolution Adaptation (C3A), which not only achieves high-rank adaptation with enhanced performance but also excels in both computational power and memory utilization. Extensive experiments demonstrate that C3A consistently outperforms LoRA and its variants across various fine-tuning tasks.
1 Introduction
In recent years, Large Foundation Models (LFMs) have surged in popularity due to their remarkable performance across various tasks in natural language processing Brown et al. (2020); Touvron et al. (2023), computer vision Radford et al. (2021); Kirillov et al. (2023), and other domains Li et al. (2024). Characterized by their vast number of parameters and substantial computational demands, these models have set new benchmarks in accuracy and versatility. However, their size and complexity pose significant challenges for efficient fine-tuning, particularly in resource-constrained environments Malladi et al. (2023); Zhang et al. (2024). To address this issue, Parameter-efficient fine-tuning (PEFT) techniques Mangrulkar et al. (2022), such as Low-Rank Adaptation (LoRA) Hu et al. (2021), have emerged as practical solutions.
LoRA reduces the number of trainable parameters by employing low-rank matrices to approximate weight changes, making fine-tuning more feasible and accessible without compromising the model’s performance. Concretely, LoRA can be expressed in math:
$\mathbf{Wx}=(\mathbf{W}_{0}+\Delta\mathbf{W})\mathbf{x}=\mathbf{W}_{0}\mathbf{x}+\mathbf{B(Ax)},$
where $\mathbf{W},\mathbf{W}_{0},\Delta\mathbf{W}\in\mathbb{R}^{d\times d}$ are weight matrices, $\mathbf{B}\in\mathbb{R}^{d\times r},\mathbf{A}\in\mathbb{R}^{r\times d}$ are low-rank matrices to construct $\Delta\mathbf{W}$ and $\mathbf{x}\in\mathbb{R}^{d}$ is the activations. The number of trainable parameters is $2rd$, prompting the choice $r\ll d$ (e.g., $r=8$ for $d=1024$) for achieving high parameter efficiency. However, as discussed in Zeng & Lee (2023), LoRA’s capacity to represent a target model is strictly limited by $r$. To address this trade-off between performance and efficiency, Kopiczko et al. (2023) proposed Vector Random Matrix Adaptation (VeRA), which achieves competitive performance with significantly fewer trainable parameters through fixed random-matrix projections. Despite its low parameter count, VeRA requires substantial computational power and memory due to the large size of the random matrices used in projection. This raises the following open research question in the context of PEFT:
How can high-rank adaptation be achieved without incurring significant costs in terms of computational power, memory usage, and trainable parameters?
As discussed in Dosovitskiy et al. (2020), dense linear layers lack strong inductive biases, limiting transformers’ ability to compete with Convolutional Neural Networks (CNNs) when data is scarce. Given the constraints of a small training set for the downstream task, we posit that incorporating a stronger inductive bias could enhance adaptation. In light of this, we propose Circular Convolution Adaptation (C3A). Circular convolution is famous in signal processing Li et al. (2020) and cryptography Dworkin et al. (2001). This operator can be equivalently represented as multiplication by a circulant matrix, offering flexibility in rank unconstrained by the number of trainable parameters. Moreover, leveraging the Fast Fourier Transform (FFT), C3A ensures better computational and memory efficiency than directly multiplying a delta matrix Bamieh (2018), rendering it competitive with LoRA from the efficiency aspect. We summarize our contribution as below:
① We introduce C3A, a novel approach for PEFT using convolution. This method leverages circular convolution kernels and their equivalent circulant matrices to provide a flexible rank, which is free of linear constraint by the number of trainable parameters, for the delta matrix.
② Leveraging the elegant diagonalization of the circulant matrix, we implement both the forward pass and backpropagation using FFT. With the incorporation of FFT, the computation and memory efficiency of C3A excels. C3A strikes a unique balance between performance an efficiency.
③ To offer greater flexibility in controlling the number of trainable parameters, we extend C3A by incorporating block-circular convolution, which results in block-circulant matrices. This extension allows C3A to achieve fully customizable parameter counts as well as adaptable rank configurations.
④ We validate C3A through comprehensive fine-tuning experiments across diverse tasks including natural language understanding, instruction tuning, image classification, and text-to-image generation. Our experiments demonstrate C3A’s outstanding accuracy and memory merits compared to existing state-of-the-art methods.
2 Related Work
2.1 Parameter-Efficient Fine-Tuning
Research on Parameter-Efficient Fine-Tuning (PEFT) has generally progressed along three main directions. The first direction involves partially updating the pre-trained neural network (e.g., the body Oh et al. (2020) or the biases Zaken et al. (2021)). Initial methods relied on hand-crafted heuristics Han et al. (2015); Oh et al. (2020) to identify which parameters are crucial and should be fine-tuned. More advanced approaches employ optimization techniques Guo et al. (2020); Xu et al. (2021); Fu et al. (2023). For example, Guo et al. Guo et al. (2020) reformulated the discrete optimization problem into a continuous one using Bernoulli masks and the Gumbel-softmax approximation.
The second direction emerged to maintain the integrity of the pre-trained model and enable a high degree of parameter sharing through adapter-based methods He et al. (2021); Rebuffi et al. (2017); Rücklé et al. (2020); Liu et al. (2022). These works focus on integrating additional modules, termed adapters, to fit the downstream task, effectively decoupling the pre-trained model parameters from those specific to the downstream task. Prompt Tuning Brown et al. (2020); Gao et al. (2020); Chen et al. (2023) and Prefix Tuning Li & Liang (2021) also fall into this category, despite ignoring potential semantic meanings.
The final direction is characterized by delta-weight-based methods, such as Low-Rank Adaptation (LoRA) Hu et al. (2021) and Orthogonal Fine-tuning (OFT) Qiu et al. (2023). These methods bridge the gap between the pre-trained model and the downstream task by adaptive delta weights, which is learned and stored separately while used in combination with the pre-trained weights. This unique design enables disentanglement and parameter sharing while preserving the ability of integrating without additional inference cost. LoRA models the delta-weights by an additive matrix while OFT does it by a multiplicative one. To further improve either parameter efficiency or performance, many variants has been proposed for both of the methods Kopiczko et al. (2023); Liu et al. (2024; 2023); Yuan et al. (2024). However, these methods can hardly achieve high parameter efficiency and performance without incurring heavy computation and memory usage.
2.2 Circular Convolution
Circular convolution has been extensively studied in signal processing Rabiner et al. (1978); McGillem & Cooper (1984); Li et al. (2020) and cryptography Dworkin et al. (2001); Gong et al. (2024). Owing to its computational advantages, circular convolution has also been explored in machine learning for generating long embeddings of high-dimensional data Yu et al. (2014) and compressing heavily parameterized layers Cheng et al. (2015); Ding et al. (2017). Remarkably, it achieves these efficiencies without significant performance degradation, which makes it a promising technique for fine-tuning applications.
Despite its success in small neural networks such as LeNet Cheng et al. (2015), circular convolution has not demonstrated lossless performance in modern large foundational models (LFMs) or even in their base architecture, the transformer. This limitation may be attributed to the conflict between its high intrinsic bias (i.e., the circulant pattern) and the vast amount of data required for training LFMs. Conversely, when fine-tuning LFMs, it is often impractical to collect as much data as needed for training from scratch. In such scenarios, the intrinsic bias of circular convolution could potentially serve as a regularization mechanism, thereby benefiting the fine-tuning process.
3 Method
In this section, we present C3A (see an overview in Figure 1), a novel PEFT approach based on the circular convolution. C3A follows LoRA’s setting of learning an additive linear operation over the original dense linear transformation. However, instead of using low-rank decomposition and the matrix multiplication operator, C3A resorts to circular convolution as this additive linear operation. Section 3.1 introduces the notations we use. Section 3.2 discusses the circular convolution operator, its equivalent circulant matrix, and its calculation in the frequency domain. Section 3.3 details an efficient method for backpropagation. Section 3.4 describes block-wise convolution for controlling the number of trainable parameters. Finally, Section 3.5 analyzes the computational complexity.
Figure 1: Overview of LoRA (A) and our C3A (B,C) method. In LoRA, only low-rank matrices $\mathbf{A}$ and $\mathbf{B}$ are trained and the delta weight is represented by their product (i.e., $\Delta\mathbf{W}=\mathbf{BA}$). The total trainable parameter number is $2rd$, which is assosiated with the rank of the delta weight. In C3A, circular convolution kernels $\mathbf{w}$ are tuned to adapt to the downstream task and the delta weight is represented by the (block-)circular matrix they construct (i.e., $\Delta\mathbf{W}=\mathcal{C}_{b}(\Delta\mathbf{w})$). The total trainable parameter number is $bd$, which disentangles with the rank of the delta weight.
3.1 Notations
We denote circular convolution operator as $\star$. The adapted weight matrix, the original weight matrix, and the delta matrix are denoted by $\mathbf{W}$, $\mathbf{W}_{0}$, and $\Delta\mathbf{W}$, respectively. The activation vector of the previous layer is denoted by $\mathbf{x}$. The post-transformation vector is $\mathbf{z}$, where $\mathbf{z}=\mathbf{W}\mathbf{x}$, and the incremental part is denoted by $\Delta\mathbf{z}$, where $\Delta\mathbf{z}=\Delta\mathbf{W}\mathbf{x}$. The length of the activation vector is $d$. The matrices $\mathbf{A}$ and $\mathbf{B}$ are low-rank matrices introduced by LoRA to represent $\Delta\mathbf{W}$, with $r$ being their rank. $r_{v}$ is used to specify the random projection matrix rank used in VeRA. The circular convolution kernel of C3A is denoted by $\Delta\mathbf{w}$. The loss function is represented by $\mathcal{L}$. The Fast Fourier Transform and its inverse are denoted by $\mathrm{FFT}$ and $\mathrm{iFFT}$, respectively. The Hadamard product is denoted by $\circ$. In terms of dimensions, we have $\mathbf{W},\mathbf{W}_{0},\Delta\mathbf{W}\in\mathbb{R}^{d\times d}$, and $\mathbf{x},\mathbf{z},\Delta\mathbf{z},\Delta\mathbf{w}\in\mathbb{R}^{d}$. The matrices $\mathbf{B}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times d}$ are controlled by the rank $r$.
3.2 Circular Convolution
The circular convolution operator is defined as $\Delta\mathbf{z}=\Delta\mathbf{w}\star\mathbf{x}=\mathcal{C}(\Delta\mathbf{w})\mathbf{x}$, where $\mathcal{C}(\Delta\mathbf{w})$ is the circulant matrix of $\Delta\mathbf{w}$. The first row of $\mathcal{C}(\Delta\mathbf{w})$ is $\Delta\mathbf{w}$ and each row equals the row above it periodically shifted to right by one element. Concretely,
$\mathcal{C}(\Delta\mathbf{w})=\begin{bmatrix}\Delta w_{1}&\Delta w_{2}&\cdots&\Delta w_{d-1}&\Delta w_{d}\\ \Delta w_{d}&\Delta w_{1}&\cdots&\Delta w_{d-2}&\Delta w_{d-1}\\ \cdots&\cdots&\cdots&\cdots&\cdots\\ \Delta w_{3}&\Delta w_{4}&\cdots&\Delta w_{1}&\Delta w_{2}\\ \Delta w_{2}&\Delta w_{3}&\cdots&\Delta w_{d}&\Delta w_{1}\\ \end{bmatrix}\in\mathbb{R}^{d\times d}.$
Theoretically, the rank of $\mathcal{C}(\Delta\mathbf{w})$ is $d-\mathrm{Deg}(\mathrm{gcd}(f(x),x^{d}-1))$ Ingleton (1956), where $\mathrm{Deg}(\cdot)$ represents the degree of a polynomial, $f(x)$ is the associated polynomial of $\Delta\mathbf{w}$ (i.e., $f(x)=\sum_{i=1}^{d}\Delta w_{i}x^{i-1}$), and $\mathrm{gcd}(\cdot)$ means the greatest common divisor. By learning $\Delta\mathbf{w}$ in the $\mathbb{R}^{n}$ oracle, C3A automatically achieves dynamic rank selection that is not linearly bounded by the number of learnable parameters as LoRA.
To achieve efficiency, we leverage the beautiful circulant structure of $\mathcal{C}(\Delta\mathbf{w})$, which makes it diagonalizable by the Fourier basis ($\mathbf{F}$) . In math, it can be described as $\mathcal{C}(\Delta\mathbf{w})=\mathbf{F}\frac{\Lambda}{d}\mathbf{F}^{-1}$ Golub & Van Loan (1996), where $\Lambda$ is its eigenvalues and can be calculated by a Fourier transform of the first row (i.e., $\Lambda=\mathrm{diag}(\mathbf{F}\Delta\mathbf{w})$). Therefore, we can calculate $\Delta\mathbf{w}\star\mathbf{x}$ as
$\begin{split}\Delta\mathbf{w}\star\mathbf{x}&=\mathbf{F}\mathrm{diag}(\frac{\mathbf{F}\Delta\mathbf{w}}{d})\mathbf{F}^{-1}\mathbf{x}\\ &=\mathrm{FFT}(\mathrm{FFT}(\Delta\mathbf{w})\circ\mathrm{iFFT}(\mathbf{x})).\end{split}$ (1)
3.3 Backpropagation
To implement backpropagation efficiently, we require the analytical derivatives of the loss function $\mathcal{L}$, with respect to $\Delta\mathbf{w}$ and $\mathbf{x}$. Utilizing the chain rule, these derivatives can be expressed as:
$\frac{\partial\mathcal{L}}{\partial\mathbf{x}}=\frac{\partial\Delta\mathbf{z}}{\partial\mathbf{x}}\frac{\partial\mathcal{L}}{\partial\Delta\mathbf{z}},\qquad\frac{\partial\mathcal{L}}{\partial\Delta\mathbf{w}}=\frac{\partial\Delta\mathbf{z}}{\partial\Delta\mathbf{w}}\frac{\partial\mathcal{L}}{\partial\Delta\mathbf{z}}.$ (2)
Since $\Delta\mathbf{z}=\mathcal{C}(\Delta\mathbf{w})\mathbf{x}$, it follows that $\frac{\partial\Delta\mathbf{z}}{\partial\mathbf{x}}=\mathcal{C}(\Delta\mathbf{w})$. For $\frac{\partial\Delta\mathbf{z}}{\partial\Delta\mathbf{w}}$, we note that the circular convolution operation is commutative (i.e., $\Delta\mathbf{w}\star\mathbf{x}=\mathbf{x}\star\Delta\mathbf{w}$), thus $\frac{\partial\Delta\mathbf{z}}{\partial\Delta\mathbf{w}}=\mathcal{C}(\mathbf{x})$. Substituting these results into Equation 2, we obtain:
$\frac{\partial\mathcal{L}}{\partial\mathbf{x}}=\mathcal{C}(\Delta\mathbf{w})\frac{\partial\mathcal{L}}{\partial\Delta\mathbf{z}},\qquad\frac{\partial\mathcal{L}}{\partial\Delta\mathbf{w}}=\mathcal{C}(\mathbf{x})\frac{\partial\mathcal{L}}{\partial\Delta\mathbf{z}}.$
These expressions can also be interpreted as circular convolutions:
$\frac{\partial\mathcal{L}}{\partial\mathbf{x}}=\Delta\mathbf{w}\star\frac{\partial\mathcal{L}}{\partial\Delta\mathbf{z}},\qquad\frac{\partial\mathcal{L}}{\partial\Delta\mathbf{w}}=\mathbf{x}\star\frac{\partial\mathcal{L}}{\partial\Delta\mathbf{z}}.$
By explicitly implementing this derivative calculation using Equation 1, backpropagation can also enjoy the computational efficiency provided by the FFT algorithm.
3.4 Block-Circular Convolution
Despite the elegance and efficiency of the circular convolution operator, it is worth noting that we can not alter the number of learnable parameters due to the requirement of the convolution kernel being the same size as the activation vector. To enable this flexibility in case of difficult downstream tasks, we resort to the block-circular convolution. We can simply split the activation vector $\mathbf{x}$ (and the post-transformation vector $\mathbf{z}$) into blocks, and assign each pair of blocks with different convolution kernels. To be concrete,
$\mathbf{x}=\begin{bmatrix}\mathbf{x}_{1}&\mathbf{x}_{2}&\cdots&\mathbf{x}_{b}\end{bmatrix}\qquad\Delta\mathbf{z}=\begin{bmatrix}\Delta\mathbf{z}_{1}&\Delta\mathbf{z}_{2}&\cdots&\Delta\mathbf{z}_{b}\end{bmatrix},$
where $b$ is the block number. We will need $b^{2}$ convolution kernels to densely connect these blocks, which can be expressed as
$\Delta\mathbf{z}_{i}=\sum_{j=1}^{b}\Delta\mathbf{w}_{ij}\star\mathbf{x}_{j},i\in\{1,2,\cdots,b\},$
and can also be represented by a block-circular matrix
$\Delta\mathbf{z}=\mathcal{C}_{b}(\Delta\mathbf{w})\mathbf{x}\qquad\mathcal{C}_{b}(\Delta\mathbf{w})=\begin{bmatrix}\mathcal{C}(\Delta\mathbf{w}_{11})&\mathcal{C}(\Delta\mathbf{w}_{12})&\cdots&\mathcal{C}(\Delta\mathbf{w}_{1b})\\ \mathcal{C}(\Delta\mathbf{w}_{21})&\mathcal{C}(\Delta\mathbf{w}_{22})&\cdots&\mathcal{C}(\Delta\mathbf{w}_{2b})\\ \cdots&\cdots&\cdots&\cdots\\ \mathcal{C}(\Delta\mathbf{w}_{b1})&\mathcal{C}(\Delta\mathbf{w}_{b2})&\cdots&\mathcal{C}(\Delta\mathbf{w}_{bb})\end{bmatrix}.$
In this case, $\Delta\mathbf{w}_{ij}\in\mathbb{R}^{\frac{d}{b}}$ (it is required that $d$ is divisible by $b$) and we have $b^{2}\frac{d}{b}=bd$ learnable parameters. $b$ acts as a hyperparameter controlling the number of learnable parameters here, just like $r$ for LoRA. However, we highlight that while $r$ controls the rank of the delta matrix and learnable parameter number at the same time, $b$ only affects the latter. Disentangling these attributes enables higher adaptation flexibility and potential better results.
3.5 Complexity Analysis
We compare time compelxity and space complexity of LoRA, VeRA and C3A in Table 1. Detailed analysis follows.
Table 1: Time and space complexity comparison of LoRA, VeRA and C3A. We split the space complexity into Parameter number and Other auxiliary tensors to help better understand the differences. We highlight that in practice, to achieve similar performance, $b\leq r\ll r_{v}$.
Method Time # Parameter # Other # Total
LoRA $\mathcal{O}(rd)$ $2rd$ $r$ $2rd+r$
VeRA $\mathcal{O}(r_{v}d)$ $r_{v}+d$ $2r_{v}d+r_{v}+d$ $2(r_{v}d+r_{v}+d)$
C3A $\mathcal{O}(bd\log\frac{d}{b})$ $bd$ 00 $bd$
3.5.1 Time Complexity
LoRA incorporates low-rank matrices $\mathbf{A}$ and $\mathbf{B}$, which are sequentially multiplied with the activation vector, resulting in a time complexity of $\mathcal{O}(rd)$. Typically, $r\ll d$. Due to the $\mathcal{O}(d\log d)$ complexity of the FFT algorithm used in Equation 1, C3A achieves a time complexity of $\mathcal{O}(bd\log\frac{d}{b})$. Given that $b\ll d$, C3A is comparable to LoRA in terms of time complexity. Conversely, VeRA, despite its high-rank nature and minimal number of trainable parameters, incurs a prohibitive time complexity of $\mathcal{O}(r_{v}d)$, where $r_{v}$ can exceed $d$. Therefore, striking a balance between high rank and computational efficiency is never a low-hanging fruit.
3.5.2 Space Complexity
We analyze the space complexity of LoRA, VeRA, and C3A during training. The differences among these methods primarily arise from the trainable parameter size and the auxiliary tensors required for the forward pass and backpropagation. For LoRA, an intermediate tensor of size $r$ is introduced (i.e., $\mathbf{Ax}\in\mathbb{R}^{r}$). In contrast, VeRA necessitates two random projection matrices, with a total size of $2r_{v}d$. Additionally, VeRA introduces two intermediate tensors of sizes $d$ and $r_{v}$. Thanks to the efficient backpropagation algorithm detailed in Section 3.3, C3A does not require any additional intermediate tensors. This optimization significantly reduces its space complexity compared to LoRA and VeRA.
4 Experiment
In this section, we comprehensively evaluate C3A in both natural language processing (NLP) and computer vision (CV). For NLP, we first demonstrate the effectiveness of C3A using the GLUE benchmark with RoBERTa-Base and RoBERTa-Large models. Additionally, we apply C3A to fine-tuning the LLaMA family models, which stands as a representative of prevalent large language models (LLMs). For CV, we examine two practical scenarios: the traditional classification task using Vision Transformers (ViTs) and the more intricate text-to-image generation task based on diffusion models.
4.1 Natural Language Understanding
Baselines. We compare our C3A with several representative PEFT methods, including BitFit Zaken et al. (2021), (IA)3 Liu et al. (2022), LoRA Hu et al. (2021), VeRA Kopiczko et al. (2023), and BOFT Liu et al. (2023). BitFit selectively fine-tunes existing parameters, specifically the biases. (IA)3 is the state-of-the-art method that adds extra adapters. LoRA is a widely known PEFT method that employs low-rank decomposition to compress additive delta matrices. VeRA is a recent approach that focuses on further reducing trainable parameters for LoRA. BOFT is another innovative method in PEFT research, compressing multiplicative delta matrices using orthogonal decomposition and butterfly factorization.
Settings. We evaluate our proposed C3A on the General Language Understanding Evaluation (GLUE) benchmark Wang et al. (2018), which encompasses a wide range of natural language understanding (NLU) tasks, including single-sentence classification, similarity and paraphrase, and natural language inference. To enhance practicality, we split these datasets following the train-validation-test approach. The best-performing model is selected based on validation set performance across the fine-tuning epochs, and the reported performance corresponds to its performance on the test set. For this evaluation, we fine-tune the pre-trained RoBERTa-Base and RoBERTa-Large models Liu et al. (2019). For the unique hyperparameters of different baselines, we adopt the values suggested in the original papers (e.g., VeRA’s $r$ and BOFT’s $b$ and $m$). The number of trainable parameters excludes the classification head, as each method uses the same one. The shared hyperparameters (i.e., the learning rate for classification head and the learning rate for other trainable parameters) are found by hyperparameter search. For the memory cost, to ensure fairness and consistency, we fix the length of input data to 256 tokens and use a batchsize of 64.
Results. Results are summarized in Table 2. In general, C3Ab=1 achieves better or on-par performance compared to baseline methods with extremely small trainable parameters. When a larger number of trainable parameters is allowed, C3Ab=6 and C3Ab=8 significantly outperform the baselines. Furthermore, compared to (IA)3, LoRA, VeRA, and BOFT, despite training a similar number of or more parameters, C3A stands out for its excellent memory efficiency. The only method with better memory efficiency is BitFit, which can be seen as an upper bound since it does not introduce any new parameters.
Table 2: Performance of different PEFT methods on the GLUE benchmark. We fine-tune pre-trained RoBERTa-Base and -Large models on 6 datasets. We report the Matthew’s correlation coefficient (MCC) for CoLA, Pearson correlation coefficient (PCC) for STS-B, and accuracy (Acc.) for all remaining tasks. For all 6 datasets, a higher metric indicates better performance. “Avg.” denotes the average accuracy of each method across all datasets. The best results for each dataset are highlighted in bold. # Trainable parameters does not include the classification head since each method uses the same head. Memory Cost is measured on fixed length (i.e., $256$) data with a batchsize of $64$.
Methods # Trainable Parameters Memory Cost (GB) SST-2 MRPC CoLA QNLI RTE STS-B Avg.
Base BitFit 102K 12.60 93.30$\pm$0.30 85.80$\pm$0.21 59.21$\pm$1.74 91.96$\pm$0.18 73.07$\pm$1.34 90.18$\pm$0.17 82.25
(IA)3  [^1] K 19.86 92.98$\pm$0.34 85.86$\pm$0.59 60.49$\pm$1.09 91.56$\pm$0.17 69.10$\pm$1.18 90.06$\pm$0.21 81.67
LoRAr=8 295K 13.75 94.50$\pm$0.41 85.68$\pm$0.74 60.95$\pm$1.57 92.54$\pm$0.20 76.68$\pm$1.42 89.76$\pm$0.39 83.35
VeRAr=1024 43K 15.51 93.97$\pm$0.17 86.23$\pm$0.41 62.24$\pm$1.91 91.85$\pm$0.17 75.74$\pm$1.56 90.27$\pm$0.25 83.38
[…]
C3Ab=1 18K 12.83 93.42$\pm$0.26 86.33$\pm$0.32 61.83$\pm$0.96 91.83$\pm$0.04 76.17$\pm$1.39 90.46$\pm$0.29 83.34
C3Ab=6  [^1] K 12.72 94.20$\pm$0.16 86.67$\pm$0.54 62.48$\pm$1.20 92.32$\pm$0.25 77.18$\pm$1.41 90.16$\pm$0.42 83.84
Large BitFit 271K 30.65 95.09$\pm$0.27 88.10$\pm$0.76 65.40$\pm$0.76 94.06$\pm$0.14 82.60$\pm$1.15 91.73$\pm$0.20 86.16
(IA)3 295K 48.81 95.32$\pm$0.20 87.06$\pm$0.57 66.52$\pm$1.10 94.18$\pm$0.15 84.33$\pm$2.38 91.58$\pm$0.39 86.50
LoRAr=8 786K 34.12 95.53$\pm$0.35 86.12$\pm$3.22 65.16$\pm$0.76 93.73$\pm$0.30 83.75$\pm$0.51 91.46$\pm$0.21 85.96
VeRAr=256 61K 34.16 95.83$\pm$0.43 87.72$\pm$0.55 63.66$\pm$1.45 94.11$\pm$0.20 83.03$\pm$1.65 91.12$\pm$0.37 85.91
[…]
C3Ab=1 49K 31.83 95.78$\pm$0.05 88.02$\pm$0.62 66.59$\pm$1.20 94.22$\pm$0.25 82.89$\pm$0.67 91.86$\pm$0.14 86.56
C3Ab=8 393K 31.79 95.78$\pm$0.15 88.09$\pm$0.47 67.18$\pm$1.92 94.26$\pm$0.19 84.62$\pm$1.36 91.81$\pm$0.36 86.96
5 Conclusion
In this paper, we introduce C3A, an innovative approach to Parameter-Efficient Fine-Tuning (PEFT). Unlike LoRA, which utilizes low-rank decomposition, C3A employs circular convolution and its associated circulant matrix to model the delta weight matrix. This approach aims to decouple the rank of the delta weight matrix from the number of learnable parameters, allowing for high-rank adaptation while maintaining a limited parameter size. By leveraging the Fast Fourier Transform (FFT) in both the forward and backward passes, C3A achieves significant computational and memory efficiency. Consequently, C3A presents itself as a compelling alternative to LoRA for model fine-tuning.
Appendix A Appendix